OpenStudy (anonymous):

if y varies directly as x and z, and y = 6 when x = 9 and z = 3, find y when x = 7 and z = 2

OpenStudy (ness9630):

Here is your starting equation: y=kxz Now solve for k

OpenStudy (anonymous):

k = 2/9, so y = 28/9

OpenStudy (ness9630):

That's right, nice job :)

OpenStudy (anonymous):

i have one more that i can't figure out, here it is \[-\frac{ x^2 }{ y }=3\]

OpenStudy (ness9630):

what are doing for this one?

OpenStudy (anonymous):

finding the constant variation. is the constant variation 3?

OpenStudy (anonymous):

it can't be 3.. because x is squared. right?

OpenStudy (ness9630):

What exactly is the question?

OpenStudy (anonymous):

it just says to find the constant variation of that equation

OpenStudy (anonymous):

i know that y = kx and if i knew how to rewrite that equation in terms of k, i'd probably be able to figure it out.

OpenStudy (ness9630):

There's no constant though..

OpenStudy (anonymous):

it says -1/3, but i dont know how they got that then

OpenStudy (anonymous):

just a guess here, but multiply y to both sides: \[-x^{2} = 3y\]then divide each side by "3" \[-\frac{ 1 }{ 3 } x^2 = y\] the constant of variation would be the "-1/3" times the parabola "x^2"

OpenStudy (anonymous):

That makes sense. Thank you!